OpenStudy (anonymous):

If one side of an equilateral triangle is 12 cm find the Area?

OpenStudy (kropot72):

Do you know the formula for the area of a triangle?

OpenStudy (anonymous):

1/2 * base * height?

OpenStudy (kropot72):

Thats right. So to use the formula we need to find the vertical height.  In the drawing sin 60 = h/6 Can you find h ?

OpenStudy (anonymous):

would that be 15.20

OpenStudy (anonymous):

3 x the square root of 3?

OpenStudy (kropot72):

Sorry my drawing has a mistake. The side of the triangle is 12 cm. \[\sin 60=\frac{\sqrt{3}}{2}=\frac{h}{12}\] \[h=6\times \sqrt{3}\] Can you find the area of the triangle now?